OpenStudy (anonymous):

The half-life of radium is 1600 years. If the initial amount is q0 milligrams, then the quantity q(t) remaining after t years is given by q(t) = q02kt. Find k.

OpenStudy (anonymous):

something wrong here

OpenStudy (anonymous):

so after 1600 years the amount q0 has halved. i.e. q0/2=q0*2*k*1600, rearrange and solve for k

OpenStudy (anonymous):

maybe \[Q(t)=Q_0\times 2^{kt}\]?

OpenStudy (anonymous):

how can we solve it??????

OpenStudy (anonymous):

the half life is \(1600\) years, so the model will be \[Q=Q_0\times \left(\frac{1}{2}\right)^{\frac{t}{1600}}\]

OpenStudy (anonymous):

oh okay thanks

OpenStudy (anonymous):

since you are using a base of 2 instead of \(\frac{1}{2}\) you can write it as \[Q=Q_0\times 2^{-\frac{t}{1600}}\]

OpenStudy (anonymous):

so i guess \(k=-\frac{1}{1600}\)

OpenStudy (anonymous):

- 1/1600